### v1

# The Size-Mass relation at Rest-Frame m from JWST/NIRCam in the COSMOS-WEB and PRIMER-COSMOS fields

Abstract
We present the galaxy stellar mass - size relation in the rest-frame near-IR () and its evolution with redshift up to . Sérsic profiles are measured for galaxies with stellar masses from JWST/NIRCam F277W and F444W imaging provided by the COSMOS-WEB and PRIMER surveys, using coordinates, redshifts, colors and stellar mass estimates from the COSMOS2020 catalog. The new rest-frame near-IR effective radii are generally smaller than previously measured rest-frame optical sizes, on average by 0.14 dex, with no significant dependence on redshift. For quiescent galaxies this size offset does not depend on stellar mass, but for star-forming galaxies the offset increases from -0.1 dex at to -0.25 dex at . That is, we find that the near-IR stellar mass - size relation for star-forming galaxies is flatter in the rest-frame near-IR than in the rest-frame optical at all redshifts . The general pace of size evolution is the same in the near-IR as previously demonstrated in the optical, with slower evolution () for star-forming galaxies and faster evolution () for quiescent galaxies. Massive () star-forming galaxies evolve in size almost as fast as quiescent galaxies. Low-mass () quiescent galaxies evolve as slow as star-forming galaxies. Our main conclusion is that the size evolution narrative as it has emerged over the past two decades does not radically change when accessing with JWST the rest-frame near-IR, a better proxy of the underlying stellar mass distribution.
[…]
1 INTRODUCTION
The size of a galaxy retains information about its evolutionary history, and is partially a reflection of the properties of the host dark matter halo, most notably the virial radius and the angular momentum (fall80; dalcanton97; mo98). Galaxy sizes are therefore a key ingredient of the various well-known scaling relations such as the size-stellar mass relation (e.g. shen03; van-der-wel14; whitaker17; mowla19a; dimauro19; Nedkova21; magnelli23; shen23; ward24; cutler24) and the fundamental plane (kormendy77; djorgovski87; dressler87).
Parametric fits can give estimates of galaxies’ size at the expense of having some sensitivity to the parametric form, but with the advantage of being less sensitive to PSF smearing and noise effects than non-parametric methods (for a complete review see conselice14). Sérsic fits are the most common profiles adopted to fit galaxies. These are parameterized by flux, half-light radius (), and Sérsic index () which quantifies the concentration of the light profile.
While the Sérsic index remains relatively constant when moving from optical wavelengths to the near-IR (kelvin12; kennedy15; martorano23; yao23), the half-light radius is significantly influenced by color gradients due to stellar population variations and attenuation by dust (sandage72; peletier90; de-jong96a). Numerous low redshift studies have analyzed the wavelength-dependent behavior of (e.g., kelvin12; haussler13; pastrav13; vika13; vulcani14; kennedy15; lange15; baes20; nersesian23). The general result is that galaxy sizes are smaller at longer wavelengths, regardless of type or mass. The implied color gradients are varyingly attributed to gradients in age, metallicity, star-formation activity and dust attenuation, depending on the type and mass of the galaxy.
Spatially resolved imaging is essential to study structural evolution at large look-back time to provide stringent constraints on galaxy evolution models. Only the Hubble Space Telescope (HST) and the James Webb Space Telescope (JWST) have sufficient spatial resolution () and stability of the point spread function to accurately quantify the light profiles of such distant galaxies. HST has been extensively used over the past two decades to map the size evolution at rest-frame UV and optical wavelengths (daddi05; trujillo06b; zirm07; van-dokkum08; van-der-wel08b; newman12; van-der-wel14; shibuya15; mowla19a; mowla19b; Nedkova21; cutler22), where attenuation by dust and the outshining by young populations present significant obstacles for interpreting the observed sizes. Now that JWST provides us with rest-frame near-IR observations of high-redshift galaxies, we can estimate sizes at longer wavelengths where dust attenuation is much less severe and the outshining effect lessened, such that gradients in the dust content and stellar population properties are suppressed.
The first estimates of rest-frame near-IR galaxy sizes from JWST by suess22 found a systematic 9% reduction in the sizes at 4.4 compared to those at 1.5 (both in the observed frame) in the redshift range . The general result that galaxies are smaller in the rest-frame near-IR compared to the rest-frame optical was subsequently confirmed (ito24; ormerod24; cutler24), displaying the same behavior as for present-day galaxies (evans94; mollenhoff06; kelvin14).
In this work, we present rest-frame 1.5 effective radius measurements for galaxies with stellar masses in the redshift range that fall in the COSMOS-WEB (casey22; casey23) and PRIMER-COSMOS (dunlop21) JWST footprints. We take advantage of these new measurements to determine the rest-frame near-IR size-mass relation and its evolution with redshift for the general galaxy population, allowing for a direct comparison with the well-established results obtained in the rest-frame optical wavelength regime.
The paper is structured as follows: in Section 2 we briefly present the cataloged data used in this work along with an overview of the JWST observations. Section 3 follows with an explanation of the fitting technique adopted to retrieve galaxies’ sizes. In Section 4 we present the size-mass relation and the redshift evolution of the median size of galaxies. This is followed by a discussion in Section LABEL:sec:discussion. Finally, in Section LABEL:sec:conclusion, we summarize our findings and draw our conclusions. We adopt a flat CDM cosmology with Hkm sMpc-1, .
2 DATA AND SAMPLE SELECTION
In this section we describe the parent catalog that we used to select target galaxies, and their basic properties (redshifts, stellar masses, star-formation activity). We also describe the JWST data used to fit near-IR light profiles.
2.1 COSMOS2020
The parent sample is drawn from the COSMOS2020-CLASSIC catalog presented by weaver22. It contains multi-wavelength ( to ) observations of over 1.7 million sources in the 2 deg2 COSMOS field (scoville07; koekemoer07) from several ground-based (i.e. Subaru/HSC and VISTA) and space-based (i.e. HST and Spitzer) observatories. The catalog includes stellar population parameters obtained from broadband photometry SED fitting with the codes LePhare (arnouts02; ilbert06) and Eazy (brammer08). Here we use the redshifts and stellar population parameters from LePhare as those are preferred by the authors (weaver22). Eazy provides the corresponding rest-frame colors UV and VJ. Mixing stellar masses from LePhare with rest-frame colors from Eazy does not bias our results: the Eazy stellar mass estimates produce the same conclusions but weaver22 prefer LePhare because of its lower redshift bias fraction. Furthermore, the LePhare stellar-mass estimates are closer to those used by van-der-wel14 that we use as a comparison sample in this work (Appendix LABEL:app-sec:_Fast-LePhare).
2.2 JWST observations
A region of deg2 roughly in the middle of the deg2 COSMOS field (Fig. 1) is covered by JWST/NIRCam imaging from the COSMOS-WEB program (GO:1727, casey23). COSMOS-WEB provides imaging with F115W and F150W in the Short Wavelength Channel and with F277W and F444W in the Long Wavelength Channel. The 5 depth reached is between 27.61 and 28.18 ABmag for the filter F444W depending on the number of exposures (casey23). Currently, slightly more than half of the deg2 area has been observed, and will be used in this work.
The CANDELS-COSMOS field (grogin11; koekemoer11), consisting of a deeper but smaller field within COSMOS, was observed by JWST as part of the PRIMER program (GO:1827, dunlop21) with ten different JWST/NIRCam filters and a F444W depth up to 28.17 ABmag (cutler24) for the regions with the highest exposures.
Data from these two programs were processed with grizli (grizli) to create mosaics and weight maps with a pixel scale of /pix. The mosaics are publicly available on The DAWN JWST Archive (DJA) (valentino23). For this work we retrieve cutouts for each galaxy from the Interactive Map Interface (hausen22) available on the DJA. Focusing on the rest-frame near-IR, in this paper we only use the Long Wavelength Channel filters F277W and F444W. As described below noise properties and PSF are tested and well-understood for this dataset.
Figure 2 shows JWST/NIRCam imaging for 16 galaxies in the sample selected to be representative of the median characteristics of galaxies across the full mass and redshift range (see Sec. LABEL:sec:discussion for further discussion). JWST/NIRCam Short Wavelength channel filters F115W and F150W are used only for visualization and are not used for quantitive analysis in this work.
2.3 Galaxy selection
Since this paper aims to present near-IR sizes at rest-frame 1.5 the wavelength coverage of the F277W and F444W filters impose a redshift range of to avoid the effects of extrapolation. The rest-frame wavelength corresponding to the pivot wavelength of F277W is 1.85 at and of F444W 1.26 for .
Among the over 1.7 million sources in the COSMOS2020-CLASSIC catalog, we select those galaxies (ACSMUCLASS=1; this flag rejects stellar-like objects) in this redshift range that have a positive HST/F814W flux and that fall within the available JWST/NIRCam F277W and F444W imaging footprints (Fig. 1).
In Figure 3 we show the mass-redshift distribution of these galaxies. The adopted stellar mass limit for our sample () leaves us with galaxies.
We remove from our sample objects (15) due to differences between the identification, deblending and segmentation between the source catalog from weaver22 and our work based on the much higher resolution NIRCam imaging (see 3.1). This rejection does not bias our sample since these targets are evenly distributed across the full mass and redshift range. Moreover, mismatches are often related to nearby bright sources and/or severe blending that prevent robust profile fits in the first place. NIRCam-selected catalogs will eventually produce a clear improvement in object selection and characterization but this is beyond the scope of this work.
Size estimates may be biased due to the contributions from central point sources from Active Galactic Nuclei (AGN). Star-like objects were already removed from the sample, so that the most obvious AGN are automatically omitted. But AGN contributions in the mid-IR still have to be addressed since chang17 showed that an AGN fraction induce a radius decrease up to . We match our sample to the publicly available catalog presented in chang17, who fit MAGPHYS on the COSMOS2015 (laigle16) photometry for sources in the COSMOS field. We find, and remove from our sample, 45 galaxies with the median of the posterior distribution of the AGN fraction larger than . Since these 45 galaxies show no obvious bias in the size distribution we conclude that more moderate AGN contributions will not affect our results either.
2.4 Separating Star-Forming and Quiescent Galaxies
The COSMOS2020 SED fits from weaver22 do not include mid- or far-IR photometric data, consequently, the resulting SFRs are not robust enough to be efficiently used to divide galaxies in star-forming and quiescent. Instead, we separate star-forming and quiescent galaxies based on their rest-frame and colors.
williams09 defined selection criteria to separate the quiescent and the star-forming population from the UVJ diagram. muzzin13 adapted the exact constraints to better match their rest-frame photometry. But we find that those cuts do not work well for the current rest-frame photometry. This difference originates from two main factors: the COSMOS2020 photometry is different from the one used in muzzin13 despite the similar sample (different methodology for photometry retrieval, several new filters, updated zeropoints); the templates used for SED fitting in COSMOS2020 are different from those used by muzzin13 leading to different rest-frame colors. These two effects combined necessitate new color cuts to maximize the separation between the two galaxy types. Note that drawing these separation lines has always been a heuristic process. In Figure 4 we present the UVJ diagram and the new color cuts adopted to optimize the separation between the two types. The quiescent region is finally identified as:
The original muzzin13 color classification only leads substantially different sample separation at , due to the many quiescent galaxies with relatively blue colors.
3 Light Profile Fits
In this section we present the steps light fitting procedure with GalfitM (haussler13; vika13), a further developed version of Galfit (peng02; peng10).
3.1 Cutouts
For each galaxy we create F277W and F444W cutouts such that the number of pixels not associated with a source in the F444W segmentation map (created with SEP, bertin96; barbary16, assuming a detection threshold of 3, a minimum area of 5 pixels and a minimum contrast ratio used for object deblending of 0.02) is at least 100 times the number of pixels that are associated with object segments. In addition, we impose a minimum size of and a maximum size of . A posteriori we verify that 99%(50%) of the cutouts are at least 12(37) times larger than , which guarantees that uncertainties in background estimates propagate into uncertainties in by less than 5%(1%) (see Fig. 6 in george24). The 1% of galaxies with cutouts smaller than 12 times are large () low-mass galaxies () mostly at . A systematic variation in the size of these targets have a negligible effect () on the median stellar mass-size relation presented in §4.1.
For each cutout we use the available weight (inverse variance) image to compute an error map. Masks are created based on the segmentation map, including those sources that are more than a magnitude fainter than the target as also done in martorano23. Sources brighter than this threshold are included in the GalfitM profile fit as separate objects, along with the target itself and fitted simultaneously. SEP provides us with an initial estimate for GalfitM of the semi-major and minor axis of the galaxies as well as their center coordinates.
We find 144 galaxies to have at least 1 pixel with flux zero within a square of side 7 pixels from the center recovered with SEP. These pixels can either be bad pixels masked by the pipeline used to create the mosaics or gaps in the footprint. Since the effect on the fit produced by a masked pixel in the inner part of the galaxy can be large, we decided to remove these 144 sources from our sample.
3.2 The Point Spread Function
GalfitM generates Sérsic profiles that are convolved with the Point-Spread-Function (PSF) of the filter investigated and subsequently fitted to the image. As several works pointed out (nardiello22; weaver24; zhuang23; ji23), we also find the PSF in the mosaics to be broader than the theoretical PSFs from WebbPSF (perrin14), which is primarily the result of the data reduction process (shifting, weighting and stacking). To minimize the systematic effects that a wrongly calibrated PSF can induce, we create empirical PSFs. We visually select 50 stars across the mosaic that are not saturated and that overlap flux density distribution with the sample galaxies as the PSF might be flux dependent. Sub-pixel precisions of the centers of the stars are determined by SEP run on cutouts. The flux is normalized in each cutout to the flux measured within a circle around the SEP center. After masking any other source in the cutout, we compute the background as the median flux per pixel in an annulus with inner radius and outer radius . This median background is subtracted from all the pixels. Each cutout is then upsampled by a factor 10 and cropped to a square of centered with sub-pixel resolution. We then produce a median stack (on a pixel-by-pixel basis) of the 50 upsampled, background-subtracted cutouts. Our final empirical PSF is then produced by a normalization such that the total flux in the square matches that of the theoretical PSF (generated via WebbPSF) in the same area. Exposures from COSMOS-WEB and PRIMER-COSMOS have the same pointing position angle (differences are degree), so that the stacked stars are well-aligned.
In Figure 5 we compare our empirical PSF with the theoretical WebbPSF PSF. The large differences in enclosed light (up to % at for F277W) imply that the WebbPSF PSF cannot be used in combination with these mosaics to infer accurate structural properties of galaxies. As referred to before, these differences can be attributed to the weighting and stacking procedure of multiple exposures in the data reduction pipeline and do not in any way reflect an error in the WebbPSF models.
The variation in stellar light profiles reflects, for one, spatial variation in the PSF. The median and 16-84 percentile ranges in the radius encircling half of the stellar light profile are . We interpret the scatter among the stellar profiles as the uncertainty in the PSF width ( and in F277W F444W). These variations are comparable to the 20% spatial variation across the detector itself found by nardiello22. The and are propagated into the galaxy size uncertainties, by adding in quadrature to the formal fitting uncertainties. The PSF uncertainty dominates the random uncertainty for small objects kpc, but does not affect the median trends that are the main result of this paper.
3.3 Profile Fits with GalfitM
The convolution box used by GalfitM is set to be 200 pixels per side. Despite slowing down the computation, the large convolution box was adopted to encompass even the largest sources. We can then use the same convolution box for all the targets avoiding the introduction of systematics due to the use of different convolution boxes for different galaxies.
Sérsic profile fits are independently fitted to the F277W and F444W images, allowing the Sérsic index to span the range 0.2 - 10 with an initial guess of 2.3. The x and y coordinates of the galaxies’ centers are limited to lie within a distance of 5 pixels from the SEP-based centers. The effective radius is limited to values between 0.01 and 200 pixels with the initial value defined as the semi-major axis of the SEP ellipse. Axis ratio and position angle are left as free parameters with no constraints. The mosaics are background-subtracted, but to allow for remaining background patterns across the mosaic we adopt a constant background with GalfitM as a free parameter.
We find that the fits reach the numerical convergence for both JWST/NIRCam filters for galaxies ( of the available sample).
We find, and remove, one galaxy with a retrieved Sérsic index within from the boundaries of the parameter space suggesting the sampler got stuck there.
For 514 sources GalfitM produces a fit with . A visual inspection of these fits reveals that for the vast majority (506) the is driven by those galaxies fitted simultaneously with the target, and not by the target itself. We refit these sources, masking the nearby galaxies and we find that they converge properly, with much lower for 497 out of 506. The remainder are omitted from the final sample.
Visual inspection by MM of all fitted galaxies shows 36 mergers and 2 lensing systems that are not properly fitted and, therefore, discarded (6 mergers and the 2 lensing systems were among those fits with ). We further reject 5 galaxies that were only partially covered by the NIRCam mosaics, leaving the fitting results unreliable. Our final sample with reliable Sérsic profiles includes galaxies of which are quiescent and are star-forming.
In Figure 6 we show the GalfitM fitting results for three randomly selected galaxies. Although low redshift high-mass galaxies are bright and extended, our pipeline uses sufficiently large cutouts to enable accurate background estimates. At the same time, for low-mass high redshift galaxies we see that the SNR is sufficiently high for a precise size estimate (larger than 30 (50) for 84% (50%) of galaxies with and redshift ).
The uncertainties retrieved for Sérsic index () and effective radius are systematically underestimated by GalfitM (van-der-wel12). Following the prescriptions in van-der-wel12 and martorano23, we reassess the uncertainties: we adopt a baseline uncertainty of 0.1 dex on the Sérsic index for galaxies with SNR = 50 and on for galaxies with SNR = 20, and compute the final uncertainties as . This methodology is based on an analysis of HST/WFC3 data (van-der-wel12). The SNR of JWST/NIRCam is much higher, producing, at times, very small uncertainties. In fact, for of the galaxies the GalfitM uncertainty on () is larger than the SNR-based value, and we retain the former.
The rest-frame 1.5 and are calculated by fitting a straight line to the values of and in F277W and F444W, using the respective pivot wavelengths. This fit is repeated 100 times, sampling from the F277W and F444W and uncertainties, adopting the median values of the and polynomials at 1.5 as the estimates and half of the 16th-84th percentile range as the 1 uncertainty.
4 RESULTS
4.1 Size-Mass Distribution

### v2

@@
 Low-mass () quiescent galaxies evolve as slow as star-forming galaxies.
-Our main conclusion is that the size evolution narrative as it has emerged over the past two decades does not radically change when accessing with JWST the rest-frame near-IR, a better proxy of the underlying stellar mass distribution.
+Our main conclusion is that the size evolution narrative as it has emerged over the past two decades does not radically change when accessing the rest-frame near-IR with JWST, a better proxy of the underlying stellar mass distribution.
 […]
 1 INTRODUCTION
-The size of a galaxy retains information about its evolutionary history, and is partially a reflection of the properties of the host dark matter halo, most notably the virial radius and the angular momentum (fall80; dalcanton97; mo98).
+The size of a galaxy retains information about its evolutionary history and is partially a reflection of the properties of the host dark matter halo, most notably the virial radius and the angular momentum (fall80; dalcanton97; mo98).
 Galaxy sizes are therefore a key ingredient of the various well-known scaling relations such as the size-stellar mass relation (e.g.
-shen03; van-der-wel14; whitaker17; mowla19a; dimauro19; Nedkova21; magnelli23; shen23; ward24; cutler24) and the fundamental plane (kormendy77; djorgovski87; dressler87).
-Parametric fits can give estimates of galaxies’ size at the expense of having some sensitivity to the parametric form, but with the advantage of being less sensitive to PSF smearing and noise effects than non-parametric methods (for a complete review see conselice14).
+shen03; van-der-wel14; whitaker17; mowla19a; dimauro19; Nedkova21; magnelli23; shen23; ward23; cutler23) and the fundamental plane (kormendy77; djorgovski87; dressler87).
+Parametric fits can give estimates of galaxies’ size at the expense of having some sensitivity to the parametric form but with the advantage of being less sensitive to point-spread function (PSF) smearing and noise effects than nonparametric methods (for a complete review, see conselice14).
 Sérsic fits are the most common profiles adopted to fit galaxies.
-These are parameterized by flux, half-light radius (), and Sérsic index () which quantifies the concentration of the light profile.
+These are parameterized by flux, half-light radius (), and Sérsic index (), which quantifies the concentration of the light profile.
 While the Sérsic index remains relatively constant when moving from optical wavelengths to the near-IR (kelvin12; kennedy15; martorano23; yao23), the half-light radius is significantly influenced by color gradients due to stellar population variations and attenuation by dust (sandage72; peletier90; de-jong96a).
-Numerous low redshift studies have analyzed the wavelength-dependent behavior of (e.g., kelvin12; haussler13; pastrav13; vika13; vulcani14; kennedy15; lange15; baes20; nersesian23).
+Numerous low-redshift studies have analyzed the wavelength-dependent behavior of (e.g., kelvin12; haussler13; pastrav13; vika13; vulcani14; kennedy15; lange15; baes20; nersesian23).
 The general result is that galaxy sizes are smaller at longer wavelengths, regardless of type or mass.
-The implied color gradients are varyingly attributed to gradients in age, metallicity, star-formation activity and dust attenuation, depending on the type and mass of the galaxy.
-Spatially resolved imaging is essential to study structural evolution at large look-back time to provide stringent constraints on galaxy evolution models.
-Only the Hubble Space Telescope (HST) and the James Webb Space Telescope (JWST) have sufficient spatial resolution () and stability of the point spread function to accurately quantify the light profiles of such distant galaxies.
+The implied color gradients are varyingly attributed to gradients in age, metallicity, star-formation activity, and dust attenuation, depending on the type and mass of the galaxy.
+Spatially resolved imaging is essential to study structural evolution at large lookback time to provide stringent constraints on galaxy evolution models.
+Only the Hubble Space Telescope (HST) and the James Webb Space Telescope (JWST) have sufficient spatial resolution () and stability of the PSF to accurately quantify the light profiles of such distant galaxies.
 HST has been extensively used over the past two decades to map the size evolution at rest-frame UV and optical wavelengths (daddi05; trujillo06b; zirm07; van-dokkum08; van-der-wel08b; newman12; van-der-wel14; shibuya15; mowla19a; mowla19b; Nedkova21; cutler22), where attenuation by dust and the outshining by young populations present significant obstacles for interpreting the observed sizes.
@@
 The first estimates of rest-frame near-IR galaxy sizes from JWST by suess22 found a systematic 9% reduction in the sizes at 4.4 compared to those at 1.5 (both in the observed frame) in the redshift range .
-The general result that galaxies are smaller in the rest-frame near-IR compared to the rest-frame optical was subsequently confirmed (ito24; ormerod24; cutler24), displaying the same behavior as for present-day galaxies (evans94; mollenhoff06; kelvin14).
+The general result that galaxies are smaller in the rest-frame near-IR compared to the rest-frame optical was subsequently confirmed (ito24; ormerod24; cutler23), displaying the same behavior as for present-day galaxies (evans94; mollenhoff06; kelvin14).
 In this work, we present rest-frame 1.5 effective radius measurements for galaxies with stellar masses in the redshift range that fall in the COSMOS-WEB (casey22; casey23) and PRIMER-COSMOS (dunlop21) JWST footprints.
 We take advantage of these new measurements to determine the rest-frame near-IR size-mass relation and its evolution with redshift for the general galaxy population, allowing for a direct comparison with the well-established results obtained in the rest-frame optical wavelength regime.
-The paper is structured as follows: in Section 2 we briefly present the cataloged data used in this work along with an overview of the JWST observations.
+The paper is structured as follows.
+In Section 2 we briefly present the cataloged data used in this work along with an overview of the JWST observations.
 Section 3 follows with an explanation of the fitting technique adopted to retrieve galaxies’ sizes.
@@
 2 DATA AND SAMPLE SELECTION
-In this section we describe the parent catalog that we used to select target galaxies, and their basic properties (redshifts, stellar masses, star-formation activity).
+In this section, we describe the parent catalog that we used to select target galaxies and their basic properties (redshifts, stellar masses, star-formation activity).
 We also describe the JWST data used to fit near-IR light profiles.
@@
 The parent sample is drawn from the COSMOS2020-CLASSIC catalog presented by weaver22.
-It contains multi-wavelength ( to ) observations of over 1.7 million sources in the 2 deg2 COSMOS field (scoville07; koekemoer07) from several ground-based (i.e.
+It contains multiwavelength ( to ) observations of over 1.7 million sources in the 2 deg2 COSMOS field (scoville07; koekemoer07) from several ground-based (i.e.
 Subaru/HSC and VISTA) and space-based (i.e.
 HST and Spitzer) observatories.
-The catalog includes stellar population parameters obtained from broadband photometry SED fitting with the codes LePhare (arnouts02; ilbert06) and Eazy (brammer08).
+The catalog includes stellar population parameters obtained from broadband photometry spectral energy distribution (SED) fitting with the codes LePhare (arnouts02; ilbert06) and Eazy (brammer08).
 Here we use the redshifts and stellar population parameters from LePhare as those are preferred by the authors (weaver22).
 Eazy provides the corresponding rest-frame colors UV and VJ.
-Mixing stellar masses from LePhare with rest-frame colors from Eazy does not bias our results: the Eazy stellar mass estimates produce the same conclusions but weaver22 prefer LePhare because of its lower redshift bias fraction.
-Furthermore, the LePhare stellar-mass estimates are closer to those used by van-der-wel14 that we use as a comparison sample in this work (Appendix LABEL:app-sec:_Fast-LePhare).
+Mixing stellar masses from LePhare with rest-frame colors from Eazy does not bias our results: the Eazy stellar mass estimates produce the same conclusions, but weaver22 prefer LePhare because of its lower redshift bias fraction.
+Furthermore, the LePhare stellar mass estimates are closer to those used by van-der-wel14, which we use as a comparison sample in this work (Appendix LABEL:app-sec:_Fast-LePhare).
 2.2 JWST observations
@@
 1) is covered by JWST/NIRCam imaging from the COSMOS-WEB program (GO:1727, casey23).
-COSMOS-WEB provides imaging with F115W and F150W in the Short Wavelength Channel and with F277W and F444W in the Long Wavelength Channel.
+COSMOS-WEB provides imaging with F115W and F150W in the short-wavelength channel and with F277W and F444W in the long-wavelength channel.
 The 5 depth reached is between 27.61 and 28.18 ABmag for the filter F444W depending on the number of exposures (casey23).
-Currently, slightly more than half of the deg2 area has been observed, and will be used in this work.
-The CANDELS-COSMOS field (grogin11; koekemoer11), consisting of a deeper but smaller field within COSMOS, was observed by JWST as part of the PRIMER program (GO:1827, dunlop21) with ten different JWST/NIRCam filters and a F444W depth up to 28.17 ABmag (cutler24) for the regions with the highest exposures.
+Currently, slightly more than half of the deg2 area has been observed and will be used in this work.
+The CANDELS-COSMOS field (grogin11; koekemoer11), consisting of a deeper but smaller field within COSMOS, was observed by JWST as part of the PRIMER program (GO:1827, dunlop21) with 10 different JWST/NIRCam filters and an F444W depth up to 28.17 ABmag (cutler23) for the regions with the highest exposures.
 Data from these two programs were processed with grizli (grizli) to create mosaics and weight maps with a pixel scale of /pix.
-The mosaics are publicly available on The DAWN JWST Archive (DJA) (valentino23).
-For this work we retrieve cutouts for each galaxy from the Interactive Map Interface (hausen22) available on the DJA.
-Focusing on the rest-frame near-IR, in this paper we only use the Long Wavelength Channel filters F277W and F444W.
-As described below noise properties and PSF are tested and well-understood for this dataset.
+The mosaics are publicly available on The DAWN JWST Archive (DJA valentino23).
+For this work, we retrieve cutouts for each galaxy from the Interactive Map Interface (hausen22) available on the DJA.
+Focusing on the rest-frame near-IR, in this paper, we only use the long-wavelength channel filters F277W and F444W.
+As described below, noise properties and PSF are tested and well understood for this dataset.
 Figure 2 shows JWST/NIRCam imaging for 16 galaxies in the sample selected to be representative of the median characteristics of galaxies across the full mass and redshift range (see Sec.
 LABEL:sec:discussion for further discussion).
-JWST/NIRCam Short Wavelength channel filters F115W and F150W are used only for visualization and are not used for quantitive analysis in this work.
+JWST/NIRCam short-wavelength channel filters F115W and F150W are used only for visualization and are not used for quantitative analysis in this work.
 2.3 Galaxy selection
-Since this paper aims to present near-IR sizes at rest-frame 1.5 the wavelength coverage of the F277W and F444W filters impose a redshift range of to avoid the effects of extrapolation.
-The rest-frame wavelength corresponding to the pivot wavelength of F277W is 1.85 at and of F444W 1.26 for .
-Among the over 1.7 million sources in the COSMOS2020-CLASSIC catalog, we select those galaxies (ACSMUCLASS=1; this flag rejects stellar-like objects) in this redshift range that have a positive HST/F814W flux and that fall within the available JWST/NIRCam F277W and F444W imaging footprints (Fig.
+Since this paper aims to present near-IR sizes at rest-frame 1.5 the wavelength coverage of the F277W and F444W filters imposes a redshift range of to avoid the effects of extrapolation.
+The rest-frame wavelength corresponding to the pivot wavelength of F277W is 1.85 at and that of F444W 1.26 for .
+Among the over 1.7 million sources in the COSMOS2020-CLASSIC catalog, we select the galaxies (ACSMUCLASS=1; this flag rejects stellar-like objects) in this redshift range that have a positive HST/F814W flux and fall within the available JWST/NIRCam F277W and F444W imaging footprints (Fig.
 1).
 In Figure 3 we show the mass-redshift distribution of these galaxies. The adopted stellar mass limit for our sample () leaves us with galaxies.
-We remove from our sample objects (15) due to differences between the identification, deblending and segmentation between the source catalog from weaver22 and our work based on the much higher resolution NIRCam imaging (see 3.1).
-This rejection does not bias our sample since these targets are evenly distributed across the full mass and redshift range.
-Moreover, mismatches are often related to nearby bright sources and/or severe blending that prevent robust profile fits in the first place.
-NIRCam-selected catalogs will eventually produce a clear improvement in object selection and characterization but this is beyond the scope of this work.
+We remove from our sample 4912 objects (15) due to differences between the identification, deblending, and segmentation between the source catalog from weaver22 and our work based on the much-higher-resolution NIRCam imaging (see 3.1).
+This rejection does not bias our sample, since these targets are evenly distributed across the full mass and redshift range.
+Moreover, mismatches are often related to nearby bright sources and/or severe blending that prevents robust profile fits in the first place.
+NIRCam-selected catalogs will eventually produce a clear improvement in object selection and characterization, but this is beyond the scope of this work.
 Size estimates may be biased due to the contributions from central point sources from active galactic nuclei (AGN).
-Star-like objects were already removed from the sample, so that the most obvious AGN are automatically omitted.
-But AGN contributions in the mid-IR still have to be addressed since chang17 showed that an AGN fraction induce a radius decrease up to .
+Starlike objects were already removed from the sample, so that the most obvious AGN are automatically omitted.
+But AGN contributions in the mid-IR still have to be addressed since chang17 showed that an AGN fraction induces a radius decrease up to .
 We match our sample to the publicly available catalog presented in chang17, who fit MAGPHYS on the COSMOS2015 (laigle16) photometry for sources in the COSMOS field.
 We find, and remove from our sample, 45 galaxies with the median of the posterior distribution of the AGN fraction larger than .
-Since these 45 galaxies show no obvious bias in the size distribution we conclude that more moderate AGN contributions will not affect our results either.
+Since these 45 galaxies show no obvious bias in the size distribution, we conclude that more moderate AGN contributions will not affect our results either.
 2.4 Separating Star-Forming and Quiescent Galaxies
-The COSMOS2020 SED fits from weaver22 do not include mid- or far-IR photometric data, consequently, the resulting SFRs are not robust enough to be efficiently used to divide galaxies in star-forming and quiescent.
+The COSMOS2020 SED fits from weaver22 do not include mid- or far-IR photometric data; consequently, the resulting star formation rates are not robust enough to be efficiently used to divide galaxies in star-forming and quiescent.
 Instead, we separate star-forming and quiescent galaxies based on their rest-frame and colors.
-williams09 defined selection criteria to separate the quiescent and the star-forming population from the UVJ diagram.
+williams09 defined selection criteria to separate the quiescent and the star-forming population from the diagram.
 muzzin13 adapted the exact constraints to better match their rest-frame photometry.
 But we find that those cuts do not work well for the current rest-frame photometry.
-This difference originates from two main factors: the COSMOS2020 photometry is different from the one used in muzzin13 despite the similar sample (different methodology for photometry retrieval, several new filters, updated zeropoints); the templates used for SED fitting in COSMOS2020 are different from those used by muzzin13 leading to different rest-frame colors.
+This difference originates from two main factors: the COSMOS2020 photometry is different from the one used in muzzin13 despite the similar sample (different methodology for photometry retrieval, several new filters, updated zero-points), and the templates used for SED fitting in COSMOS2020 are different from those used by muzzin13 leading to different rest-frame colors.
 These two effects combined necessitate new color cuts to maximize the separation between the two galaxy types.
 Note that drawing these separation lines has always been a heuristic process.
-In Figure 4 we present the UVJ diagram and the new color cuts adopted to optimize the separation between the two types.
-The quiescent region is finally identified as:
+In Figure 4 we present the diagram and the new color cuts adopted to optimize the separation between the two types.
+The quiescent region is finally identified as
 The original muzzin13 color classification only leads substantially different sample separation at , due to the many quiescent galaxies with relatively blue colors.
 3 Light Profile Fits
-In this section we present the steps light fitting procedure with GalfitM (haussler13; vika13), a further developed version of Galfit (peng02; peng10).
+In this section, we present the steps of the light fitting procedure witGalfitM (haussler13; vika13), a further developed version of Galfit (peng02; peng10).
 3.1 Cutouts
-For each galaxy we create F277W and F444W cutouts such that the number of pixels not associated with a source in the F444W segmentation map (created with SEP, bertin96; barbary16, assuming a detection threshold of 3, a minimum area of 5 pixels and a minimum contrast ratio used for object deblending of 0.02) is at least 100 times the number of pixels that are associated with object segments.
+For each galaxy, we create F277W and F444W cutouts such that the number of pixels not associated with a source in the F444W segmentation map (created with SEP, assuming a detection threshold of 3, a minimum area of 5 pixels and a minimum contrast ratio used for object deblending of 0.02; bertin96; barbary16) is at least 100 times the number of pixels that are associated with object segments.
 In addition, we impose a minimum size of and a maximum size of .
-A posteriori we verify that 99%(50%) of the cutouts are at least 12(37) times larger than , which guarantees that uncertainties in background estimates propagate into uncertainties in by less than 5%(1%) (see Fig.
+A posteriori, we verify that 99%(50%) of the cutouts are at least 12(37) times larger than , which guarantees that uncertainties in background estimates propagate into uncertainties in by less than 5%(1%) (see Fig.
 6 in george24).
 The 1% of galaxies with cutouts smaller than 12 times are large () low-mass galaxies () mostly at .
-A systematic variation in the size of these targets have a negligible effect () on the median stellar mass-size relation presented in §4.1.
-For each cutout we use the available weight (inverse variance) image to compute an error map.
-Masks are created based on the segmentation map, including those sources that are more than a magnitude fainter than the target as also done in martorano23.
-Sources brighter than this threshold are included in the GalfitM profile fit as separate objects, along with the target itself and fitted simultaneously.
-SEP provides us with an initial estimate for GalfitM of the semi-major and minor axis of the galaxies as well as their center coordinates.
+A systematic variation in the size of these targets has a negligible effect () on the median stellar mass-size relation presented in §4.1.
+For each cutout, we use the available weight (inverse variance) image to compute an error map.
+Masks are created based on the segmentation map, including those sources that are more than a magnitude fainter than the target, as also done in martorano23.
+Sources brighter than this threshold are included in the GalfitM profile fit as separate objects, along with the target itself, and fitted simultaneously.
+SEP provides us with an initial estimate for GalfitM of the semi-major and minor axes of the galaxies as well as their center coordinates.
 We find 144 galaxies to have at least 1 pixel with flux zero within a square of side 7 pixels from the center recovered with SEP.
@@
 3.2 The Point Spread Function
-GalfitM generates Sérsic profiles that are convolved with the Point-Spread-Function (PSF) of the filter investigated and subsequently fitted to the image.
-As several works pointed out (nardiello22; weaver24; zhuang23; ji23), we also find the PSF in the mosaics to be broader than the theoretical PSFs from WebbPSF (perrin14), which is primarily the result of the data reduction process (shifting, weighting and stacking).
+GalfitM generates Sérsic profiles that are convolved with the PSF of the filter investigated and subsequently fitted to the image.
+As several works pointed out (nardiello22; weaver23; zhuang23; ji23), we also find the PSF in the mosaics to be broader than the theoretical PSFs from WebbPSF (perrin14), which is primarily the result of the data reduction process (shifting, weighting, and stacking).
 To minimize the systematic effects that a wrongly calibrated PSF can induce, we create empirical PSFs.
-We visually select 50 stars across the mosaic that are not saturated and that overlap flux density distribution with the sample galaxies as the PSF might be flux dependent.
-Sub-pixel precisions of the centers of the stars are determined by SEP run on cutouts.
+We visually select 50 stars across the mosaic that are not saturated and that overlap flux density distribution with the sample galaxies, as the PSF might be flux-dependent.
+Subpixel precisions of the centers of the stars are determined by SEP run on cutouts.
 The flux is normalized in each cutout to the flux measured within a circle around the SEP center.
@@
 This median background is subtracted from all the pixels.
-Each cutout is then upsampled by a factor 10 and cropped to a square of centered with sub-pixel resolution.
+Each cutout is then upsampled by a factor 10 and cropped to a square of centered with subpixel resolution.
 We then produce a median stack (on a pixel-by-pixel basis) of the 50 upsampled, background-subtracted cutouts.
 Our final empirical PSF is then produced by a normalization such that the total flux in the square matches that of the theoretical PSF (generated via WebbPSF) in the same area.
-Exposures from COSMOS-WEB and PRIMER-COSMOS have the same pointing position angle (differences are degree), so that the stacked stars are well-aligned.
+Exposures from COSMOS-WEB and PRIMER-COSMOS have the same pointing position angle (differences are ), so that the stacked stars are well aligned.
 In Figure 5 we compare our empirical PSF with the theoretical WebbPSF PSF.
@@
 Despite slowing down the computation, the large convolution box was adopted to encompass even the largest sources.
-We can then use the same convolution box for all the targets avoiding the introduction of systematics due to the use of different convolution boxes for different galaxies.
+We can then use the same convolution box for all the targets, avoiding the introduction of systematics due to the use of different convolution boxes for different galaxies.
 Sérsic profile fits are independently fitted to the F277W and F444W images, allowing the Sérsic index to span the range 0.2 - 10 with an initial guess of 2.3.
@@
 Axis ratio and position angle are left as free parameters with no constraints.
-The mosaics are background-subtracted, but to allow for remaining background patterns across the mosaic we adopt a constant background with GalfitM as a free parameter.
+The mosaics are background-subtracted, but to allow for remaining background patterns across the mosaic, we adopt a constant background with GalfitM as a free parameter.
 We find that the fits reach the numerical convergence for both JWST/NIRCam filters for galaxies ( of the available sample).
 We find, and remove, one galaxy with a retrieved Sérsic index within from the boundaries of the parameter space suggesting the sampler got stuck there.
-For 514 sources GalfitM produces a fit with .
-A visual inspection of these fits reveals that for the vast majority (506) the is driven by those galaxies fitted simultaneously with the target, and not by the target itself.
-We refit these sources, masking the nearby galaxies and we find that they converge properly, with much lower for 497 out of 506.
+For 514 sources, GalfitM produces a fit with .
+A visual inspection of these fits reveals that for the vast majority (506) the is driven by those galaxies fitted simultaneously with the target and not by the target itself.
+We refit these sources, masking the nearby galaxies, and we find that they converge properly, with much lower for 497 out of 506.
 The remainder are omitted from the final sample.
-Visual inspection by MM of all fitted galaxies shows 36 mergers and 2 lensing systems that are not properly fitted and, therefore, discarded (6 mergers and the 2 lensing systems were among those fits with ).
-We further reject 5 galaxies that were only partially covered by the NIRCam mosaics, leaving the fitting results unreliable.
+Visual inspection by M.M.
+of all fitted galaxies shows 36 mergers and two lensing systems that are not properly fitted and are, therefore, discarded (six mergers and the two lensing systems were among those fits with ).
+We further reject five galaxies that were only partially covered by the NIRCam mosaics, leaving the fitting results unreliable.
 Our final sample with reliable Sérsic profiles includes galaxies of which are quiescent and are star-forming.
 In Figure 6 we show the GalfitM fitting results for three randomly selected galaxies.
-Although low redshift high-mass galaxies are bright and extended, our pipeline uses sufficiently large cutouts to enable accurate background estimates.
-At the same time, for low-mass high redshift galaxies we see that the SNR is sufficiently high for a precise size estimate (larger than 30 (50) for 84% (50%) of galaxies with and redshift ).
+Although low-redshift high-mass galaxies are bright and extended, our pipeline uses sufficiently large cutouts to enable accurate background estimates.
+At the same time, for low-mass high-redshift galaxies, we see that the signal-to-noise ratio (SNR) is sufficiently high for a precise size estimate (larger than 30 (50) for 84% (50%) of galaxies with and redshift ).
 The uncertainties retrieved for Sérsic index () and effective radius are systematically underestimated by GalfitM (van-der-wel12).
@@
 The SNR of JWST/NIRCam is much higher, producing, at times, very small uncertainties.
-In fact, for of the galaxies the GalfitM uncertainty on () is larger than the SNR-based value, and we retain the former.
-The rest-frame 1.5 and are calculated by fitting a straight line to the values of and in F277W and F444W, using the respective pivot wavelengths.
+In fact, for of the galaxies, the GalfitM uncertainty on () is larger than the SNR-based value, and we retain the former.
+The rest-frame 1.5 and are calculated by fitting a straight line to the values of and in F277W and F444W using the respective pivot wavelengths.
 This fit is repeated 100 times, sampling from the F277W and F444W and uncertainties, adopting the median values of the and polynomials at 1.5 as the estimates and half of the 16th-84th percentile range as the 1 uncertainty.
+In Table 1 we present GalfitM outputs produced from the fitting of the NIRCam/F277W, F444W cutouts and the rest-frame effective radii and Sérsic indices computed as discussed above.
+A full version of the table is available online in the machine-readable format.
+Label | Unit | Description |
+Classic | — | COSMOS2020/CLASSIC Identifier [ID] |
+RAdeg | deg | COSMOS2020/CLASSIC Right Ascension, decimal degrees (J2000) [ALPHA_J2000] |
+DEdeg | deg | COSMOS2020/CLASSIC Declination, decimal degrees (J2000) [DELTA_2000] |
+lpzBEST | — | COSMOS2020/CLASSIC LePhare photo-z redshift [lp_zBEST] |
+loglpMassbest | [Msun] | Corrected (see text) COSMOS2020/CLASSIC LePhare BC03 log stellar mass [lp_mass_best] |
+Re-F277W | arcsec | GalfitM effective radius in filter F277W |
+e_Re-F277W | arcsec | Uncertainty in Re-F277W |
+n-F277W | — | GalfitM Sérsic index in filter F277W |
+e_n-F277W | — | Uncertainty in n-F277W |
+mag-F277W | mag | GalfitM AB magnitude in filter F277W |
+e_mag-F277W | mag | Uncertainty in mag-F277W |
+q-F277W | — | GalfitM axis ratio in filter F277W |
+e_q-F277W | — | Uncertainty in q-F277W |
+PA-F277W | deg | GalfitM position angle in filter F277W |
+e_PA-F277W | deg | Uncertainty in PA-F277W |
+Re-F444W | arcsec | GalfitM effective radius in filter F444W |
+e_Re-F444W | arcsec | Uncertainty in Re-F444W |
+n-F444W | — | GalfitM Sérsic index in filter F444W |
+e_n-F444W | — | Uncertainty in n-F444W |
+mag-F444W | mag | GalfitM AB magnitude in filter F444W |
+e_mag-F444W | mag | Uncertainty in mag-F444W |
+q-F444W | — | GalfitM axis ratio in filter F444W |
+e_q-F444W | — | Uncertainty in q-F444W |
+PA-F444W | deg | GalfitM position angle in filter F444W |
+e_PA-F444W | deg | Uncertainty in PA-F444W |
+Re-1.5um | arcsec | Effective radius at rest-frame |
+e_Re-1.5um | arcsec | Uncertainty in Re-1.5um |
+n-1.5um | — | Sérsic index at rest-frame |
+e_n-1.5um | — | Uncertainty in n-1.5um |
+Note. — The
+original COSMOS2020/CLASSIC labels are given in square brackets in the
+description of the columns taken or derived from that catalog.
+Table 1 is published in its entirety in
+the machine-readable format. A description of the columns is provided here
+for guidance regarding its form and content.
 4 RESULTS